OpenStudy (anonymous):

The length of an arc of a circle is 12cm. The corresponding sector area is 108cm^2. find the radius of the circle?

OpenStudy (anonymous):

Need help plz

OpenStudy (anonymous):

area of sector = radius x arc length 108 = radius x 12 radius = 108/12 radius = 9

OpenStudy (anonymous):

But isn't sector area equation is. =sector area =theta \360 x2pie x r

OpenStudy (anonymous):

that is also true but in this case you can only use this one..

OpenStudy (anonymous):

Ok thanx

OpenStudy (anonymous):

But answer in the book is 18 cm

OpenStudy (anonymous):

oh oh oh @Shaaz14 i put the wrong formula the correct one is area = 1/2 x radius x arc length 108 = 1/2 x radius x 12 radius is therefore 18

OpenStudy (anonymous):

One more is that there saying also find the angle sub tended at the centre of the circle by the arc

OpenStudy (anonymous):

yeah so you can find the angle using the formula area = 1/2 x (radius)^2 x angle 108 = 1/2 x (18)^2 x angle angle at centre = 0.6radians

OpenStudy (anonymous):

The answer is coming 38.2*

OpenStudy (anonymous):

In the book

OpenStudy (anonymous):

my answer is in degrees u can change it to degrrees 0.6 radians = 0.6/pi x 180= 38.2 degrees

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

pi radians = 180 degrees taking this ratio u can convert 0.6 radians to degrees